# JARVIS-VLA: Post-Training Large-Scale Vision Language Models to Play Visual Games with Keyboards and Mouse

### v1

@@ D.1 World Knowledge Dataset
Example of World Knowledge Dataset Question: "What are the different items that can be made using dark oak planks in Minecraft?" Answer: "In Minecraft, you can use dark oak planks to craft various items and blocks. Here are some of the items and blocks you can make with dark oak planks: 1. Dark Oak Wood Slabs: Used for flooring, roofs, and other building elements. 2. Dark Oak Wood Stairs: Ideal for creating inclined pathways or staircases. 3. Dark Oak Wood Fence: Useful for creating boundaries or enclosures around structures. 4. Dark Oak Wood Fence Gate: Allows for openings in your fences for easy access. 5. Dark Oak Wood Door: Adds entrances to your buildings. 6. Dark Oak Wood Trapdoor: Can serve as hidden entrances or functional hatches. 7. Dark Oak Wood Button: Can be placed on surfaces and used to activate various redstone mechanisms. 8. Dark Oak Wood Pressure Plate: Activates redstone contraptions when stepped on. 9. Dark Oak Wood Sign: Useful for displaying text and providing information in your world. 10. Dark Oak Wood Bed: Used to craft the frame of a bed, combined with wool blocks for a complete bed. 11. Dark Oak Wood Boat: Useful for water transportation. 12. Dark Oak Wood Chest: Can be used to craft storage chests. 13. Dark Oak Wood Bookshelf: Decorative and enhances enchantment tables. 14. Petrified Oak Slab: Made from dark oak logs submerged in water and then crafted into planks and slabs. These items enhance functionality and aesthetics in your Minecraft builds."   Question: "How can players build and use automatic farms in Minecraft, and what are some efficient designs?" Answer: "In Minecraft, automatic farms are structures designed to automatically harvest resources without requiring constant player intervention. These farms can be extremely useful for gathering various resources efficiently. Here, I’ll provide a general overview of how to build and use automatic farms, along with examples of efficient designs for different types of farms: 1. Crop Farms: Automated crop farms are used to grow and harvest crops like wheat, carrots, potatoes, and more. Efficient design: - Create a water-based system to hydrate the farmland. A water source in the middle with canals leading to the farmland is a common approach. - Use a dispenser to plant seeds or crops onto the farmland. - Attach a redstone clock to the dispenser to periodically activate it. - Collect the harvested items using hoppers and chests. 2. Mob Farms: These farms are used to automatically collect drops from hostile mobs, such as zombies, skeletons, and creepers. Efficient design: - Build a dark room where mobs will spawn naturally. Use trapdoors, pressure plates, or a combination to restrict mob movement. - Use water currents to funnel mobs into a kill chamber. - Create a killing mechanism using either lava blades, suffocation, or other methods. - Use water currents again to collect drops and funnel them into hoppers for collection. 3. Animal Farms: Automated animal farms can help you gather resources from passive mobs like cows, pigs, and chickens. Efficient design: - Create separate pens for each type of animal. - Use a breeding mechanism to maintain a stable population. Dispensers with food (like wheat for cows) can automatically breed animals. - Include a killing mechanism for excess animals, providing a steady supply of drops. - Use water currents and hoppers to collect drops into chests. 4. Tree Farms: Automated tree farms allow you to grow and harvest trees without manual effort. Efficient design: - Plant saplings in rows using dispensers and bone meal. - Use a mechanism to cut down the trees. This can be achieved using flying machines or TNT-based contraptions. - Use water currents to collect the dropped wood and saplings into hoppers and chests. 5. Fishing Farms: Automated fishing farms simulate fishing to gather loot and experience points. Efficient design: - Build a platform with water sources, a fishing rod dispenser, and a mechanism to reel in the fishing rod. - Create a hopper system to collect items caught from the fishing process. Remember that the efficiency of these designs can vary depending on your Minecraft version and specific game settings. Additionally, always consider ethical and environmental factors when designing and using automated farms in your gameplay."
@@ D.2 Vision Language Alignment Dataset
For the construction of Vision Language Alignment Dataset, we employed a multi-step approach to gather and process data. Initially, we sourced images from a variety of online platforms, including YouTube gameplay videos, contractor-provided screenshots [4], and other internet resources. We carefully selected 35,000 keyframes from these sources based on criteria such as brightness and visual complexity to ensure the quality of the dataset. Subsequently, we utilized advanced Vision-Language Models—such as GPT-4o [35], Claude 3.5 Sonnet [2], and Molmo [16]—to generate contextual image question-answer pairs, critical in creating a rich, semantically detailed dataset that bridges visual content and linguistic annotations. To enhance the reliability of the generated data, we implemented a robust query-validation pipeline. This pipeline employed Llama-3.1-72B [20] to systematically filter out ambiguous questions and validate the accuracy of the generated answers. Additionally, it included a validation step to ensure that the question-answer pairs were relevant to the associated images. To further diversify our dataset, we specifically allocated nearly half of the vision-question-answer pairs for caption generation, employing the advanced capabilities of GPT-4o. Through the implementation of our comprehensive pipeline, we successfully generated a Vision Language Alignment Dataset comprising 15,000 captions and 20,000 visual question answering dataset.
Example of Vision Language Alignment Datasets [Uncaptioned image] Question: "What tool is the player holding?" Answer: "The player is holding a pickaxe with an enchantment glow, likely a diamond pickaxe due to its color."   Question: "How many hearts of health does the player have?"" Answer: "The player has 9 hearts of health visible, indicating they are missing 1 heart."   Question: "What type of block is being mined by the player?" Answer: "The player is mining a stone block."   Question: "What color is the liquid at the top of the image?" Answer: "The liquid at the top of the image is orange, indicating it is lava."   Question: "How many inventory slots does the player have in total (visible on the hotbar)?" Answer: "The player has 9 inventory slots visible on the hotbar."
@@ D.3 Visual Grounding dataset
In 3D environment, we sample images from the contractor data. We initially identify objects that are interacting with the player at a specific moment in the video. Then implementing Backward Trajectory Relabeling [10] by sam2 [37] to accurately trace the positions of objects up to $t$ frames backward, which tend to be in more random positions.
@@ E.1 World Knowledge Evaluation
Enchant 3 What happens if the player put on an item with Curse of Binding enchant in Minecraft? It cannot be removed until the item breaks or the player dies.
@@ E.2 Vision Understanding Evaluation
Vision Understanding Evaluation Examples [Uncaptioned image] Question: What time of day is it in the game? GPT-4o: It is difficult to determine the exact time of day in the game due to the rainy weather, which can create a darkened sky. However, based on the darkness of the surroundings, it appears to be night or close to night. Rain in Minecraft makes it harder to see the exact position of the sun or the moon, but the overall dim lighting suggests it is either late evening or night time. Qwen2-VL: It is daytime in the game. The bright light and the visible sky indicate that it is daytime. JARVIS-VLA-Qwen2-VL: The time of day in the game is night. This can be determined by the dark sky and the presence of rain in the screenshot. In Minecraft, night begins when the sun sets and the moon rises. During nighttime, the sky darkens, and if it’s raining, you’ll see rain particles falling from the sky. Additionally, night is a more dangerous time in Minecraft as hostile mobs spawn more frequently.
@@ E.4 Evaluation Metric and Result
We designed a customized evaluation method to assess the performance of models in answering the questions. For World Knowledge Questions and Visual Understanding Questions, we explore the utilization of LLMs as judges. We selected GPT-4o [35], a state-of-the-art LLM to serve as the judge. The judge model first reviews the responses and compares them to a set of expertly crafted reference answers. Subsequently, the judge assigns a score of correct or incorrect. For visual grounding tasks, we directly score the responses of the evaluated model based on a rule-based approach. Below are the performances of some models we are interested in under our benchmark: Table 5.
@@
GPT-4o [1] - 96.6 1 76.7 1 - -
GPT-4o-mini [1] - 75.9 2 62.8 4 - -
Llava-Next [27] 8B 19.0 8 41.9 10 - -
Molmo-d-0924 [16] 7B 12.1 10 58.1 5 24.8 3
Llama-3.2 [33] 11B 20.7 7 44.2 9 - -
Qwen2-VL [43] 7B 17.3 9 46.5 7 16.6 5
@@
Table 5: We compared the performance of various VLMs using our benchmark, including commercial large models (GPT-4 and GPT-4-mini [35]), open-source models (Llava-Next [27], Molmo-d-0924 [16], Llama-3.2 [33], and Qwen2-VL [43]), as well as JARVIS-VLA. The results demonstrate that our method significantly enhances the core capabilities of these models, although there remains a gap when compared to state-of-the-art models.
@@ Appendix F Ablation with different Pre-trained VLMs
We compare two models, Llava-Next [27] and Qwen2-VL [43], which utilize different pretraining datasets and image processing techniques. Their raw VLM performances and post-training results on various auxiliary tasks, along with downstream imitation learning outcomes, are presented in Table 6.
@@
Methods Models World Knowledge Visual Alignment Spatial Grounding [Uncaptioned image] [Uncaptioned image]
Raw Llava-Next-8B 18.9 41.8 - 26.7 10.0
ActVLP Llava-Next-8B 55.8 60.3 - 53.3 16.6
Raw Qwen2-VL-7B 17.3 46.5 16.6 83.3 0.0
ActVLP Qwen2-VL-7B 70.7 76.7 88.0 86.7 83.3
Table 6: Ablation experiments on base model and model structure. We adopt ActVLP on Llava-Next-8B [27] and Qwen2-VL-7B [43] to validate the robustness across different base vision language models.

### v2

@@ D.1 World Knowledge Dataset
Example of World Knowledge Dataset Question: "What are the different items that can be made using dark oak planks in Minecraft?" Answer: "In Minecraft, you can use dark oak planks to craft various items and blocks. Here are some of the items and blocks you can make with dark oak planks: 1. Dark Oak Wood Slabs: Used for flooring, roofs, and other building elements. 2. Dark Oak Wood Stairs: Ideal for creating inclined pathways or staircases. 3. Dark Oak Wood Fence: Useful for creating boundaries or enclosures around structures. 4. Dark Oak Wood Fence Gate: Allows for openings in your fences for easy access. 5. Dark Oak Wood Door: Adds entrances to your buildings. 6. Dark Oak Wood Trapdoor: Can serve as hidden entrances or functional hatches. 7. Dark Oak Wood Button: Can be placed on surfaces and used to activate various redstone mechanisms. 8. Dark Oak Wood Pressure Plate: Activates redstone contraptions when stepped on. 9. Dark Oak Wood Sign: Useful for displaying text and providing information in your world. 10. Dark Oak Wood Bed: Used to craft the frame of a bed, combined with wool blocks for a complete bed. 11. Dark Oak Wood Boat: Useful for water transportation. 12. Dark Oak Wood Chest: Can be used to craft storage chests. 13. Dark Oak Wood Bookshelf: Decorative and enhances enchantment tables. 14. Petrified Oak Slab: Made from dark oak logs submerged in water and then crafted into planks and slabs. These items enhance functionality and aesthetics in your Minecraft builds."  Question: "How can players build and use automatic farms in Minecraft, and what are some efficient designs?" Answer: "In Minecraft, automatic farms are structures designed to automatically harvest resources without requiring constant player intervention. These farms can be extremely useful for gathering various resources efficiently. Here, I’$l$ provide a general overview of how to build and use automatic farms, along with examples of efficient designs for different types of farms: 1. Crop Farms: Automated crop farms are used to grow and harvest crops like wheat, carrots, potatoes, and more. Efficient design: - Create a water-based system to hydrate the farmland. A water source in the middle with canals leading to the farmland is a common approach. - Use a dispenser to plant seeds or crops onto the farmland. - Attach a redstone clock to the dispenser to periodically activate it. - Collect the harvested items using hoppers and chests. 2. Mob Farms: These farms are used to automatically collect drops from hostile mobs, such as zombies, skeletons, and creepers. Efficient design: - Build a dark room where mobs will spawn naturally. Use trapdoors, pressure plates, or a combination to restrict mob movement. - Use water currents to funnel mobs into a kill chamber. - Create a killing mechanism using either lava blades, suffocation, or other methods. - Use water currents again to collect drops and funnel them into hoppers for collection. 3. Animal Farms: Automated animal farms can help you gather resources from passive mobs like cows, pigs, and chickens. Efficient design: - Create separate pens for each type of animal. - Use a breeding mechanism to maintain a stable population. Dispensers with food (like wheat for cows) can automatically breed animals. - Include a killing mechanism for excess animals, providing a steady supply of drops. - Use water currents and hoppers to collect drops into chests. 4. Tree Farms: Automated tree farms allow you to grow and harvest trees without manual effort. Efficient design: - Plant saplings in rows using dispensers and bone meal. - Use a mechanism to cut down the trees. This can be achieved using flying machines or TNT-based contraptions. - Use water currents to collect the dropped wood and saplings into hoppers and chests. 5. Fishing Farms: Automated fishing farms simulate fishing to gather loot and experience points. Efficient design: - Build a platform with water sources, a fishing rod dispenser, and a mechanism to reel in the fishing rod. - Create a hopper system to collect items caught from the fishing process. Remember that the efficiency of these designs can vary depending on your Minecraft version and specific game settings. Additionally, always consider ethical and environmental factors when designing and using automated farms in your gameplay."
@@ D.2 Vision Language Alignment Dataset
For the construction of Vision Language Alignment Dataset, we employed a multi-step approach to gather and process data. Initially, we sourced images from a variety of online platforms, including YouTube gameplay videos, contractor-provided screenshots (Baker et al., 2022), and other internet resources. We carefully selected 35,000 keyframes from these sources based on criteria such as brightness and visual complexity to ensure the quality of the dataset. Subsequently, we utilized advanced Vision-Language Models—such as GPT-4o (OpenAI, 2023), Claude 3.5 Sonnet (Anthropic, 2024), and Molmo (Deitke et al., 2024)—to generate contextual image question-answer pairs, critical in creating a rich, semantically detailed dataset that bridges visual content and linguistic annotations. To enhance the reliability of the generated data, we implemented a robust query-validation pipeline. This pipeline employed Llama-3.1-70B (Dubey et al., 2024) to systematically filter out ambiguous questions and validate the accuracy of the generated answers. Additionally, it included a validation step to ensure that the question-answer pairs were relevant to the associated images. To further diversify our dataset, we specifically allocated nearly half of the vision-question-answer pairs for caption generation, employing the advanced capabilities of GPT-4o. Through the implementation of our comprehensive pipeline, we successfully generated a Vision Language Alignment Dataset comprising 15,000 captions and 20,000 visual question answering dataset.
Example of Vision Language Alignment Datasets [Uncaptioned image]
Question: "What tool is the player holding?" Answer: "The player is holding a pickaxe with an enchantment glow, likely a diamond pickaxe due to its color."  Question: "What type of block is being mined by the player?" Answer: "The player is mining a stone block."  Question: "What color is the liquid at the top of the image?" Answer: "The liquid at the top of the image is orange, indicating it is lava."  Question: "How many inventory slots does the player have in total (visible on the hotbar)?" Answer: "The player has 9 inventory slots visible on the hotbar."
@@ D.3 Visual Grounding dataset
In 3D environment, we sample images from the contractor data. We initially identify objects that are interacting with the player at a specific moment in the video. Then implementing Backward Trajectory Relabeling (Cai et al., 2024a) by sam2 (Ravi et al., 2024) to accurately trace the positions of objects up to $t$ frames backward, which tend to be in more random positions.
@@ E.1 World Knowledge Evaluation
Enchant 3 What happens if the player puts on an item with the Curse of Binding enchantment in Minecraft? It cannot be removed until the item breaks or the player dies.
@@ E.2 Vision Understanding Evaluation
Vision Understanding Evaluation Examples [Uncaptioned image] Question: What time of day is it in the game? GPT-4o: It is difficult to determine the exact time of day in the game due to the rainy weather, which can create a darkened sky. However, based on the darkness of the surroundings, it appears to be night or close to night. Rain in Minecraft makes it harder to see the exact position of the sun or the moon, but the overall dim lighting suggests it is either late evening or night time. Qwen2-VL: It is daytime in the game. The bright light and the visible sky indicate that it is daytime. JARVIS-VLA-Qwen2-VL: The time of day in the game is night. This can be determined by the dark sky and the presence of rain in the screenshot. In Minecraft, night begins when the sun sets and the moon rises. During nighttime, the sky darkens, and if it’s raining, you’$l$ see rain particles falling from the sky. Additionally, night is a more dangerous time in Minecraft as hostile mobs spawn more frequently.
@@ E.4 Evaluation Metric and Result
We designed a customized evaluation method to assess the performance of models in answering the questions. For World Knowledge Questions and Visual Understanding Questions, we explore the utilization of LLMs as judges. We selected GPT-4o (OpenAI, 2023), a state-of-the-art LLM to serve as the judge. The judge model first reviews the responses and compares them to a set of expertly crafted reference answers. Subsequently, the judge assigns a score of correct or incorrect. For visual grounding tasks, we directly score the responses of the evaluated model based on a rule-based approach. Below are the performances of some models we are interested in under our benchmark: Table 8.
@@
GPT-4o (Achiam et al., 2023) - 96.6 1 76.7 1 - -
GPT-4o-mini (Achiam et al., 2023) - 75.9 2 62.8 4 - -
Llava-Next (Li et al., 2024a) 8B 19.0 8 41.9 10 - -
Molmo-d-0924 (Deitke et al., 2024) 7B 12.1 10 58.1 5 24.8 3
Llama-3.2 (Meta, 2024) 11B 20.7 7 44.2 9 - -
Qwen2-VL (Wang et al., 2024b) 7B 17.3 9 46.5 7 16.6 5
@@
Table 8: We compared the performance of various VLMs using our benchmark, including commercial large models (GPT-4 and GPT-4-mini (OpenAI, 2023)), open-source models (Llava-Next (Li et al., 2024a), Molmo-d-0924 (Deitke et al., 2024), Llama-3.2 (Meta, 2024), and Qwen2-VL (Wang et al., 2024b)), as well as JARVIS-VLA. The results demonstrate that our method significantly enhances the core capabilities of these models, although there remains a gap when compared to state-of-the-art models.
@@ Appendix F Ablation with different Pre-trained VLMs
We compare two models, Llava-Next (Li et al., 2024a) and Qwen2-VL (Wang et al., 2024b), which utilize different pretraining datasets and image processing techniques. Their raw VLM performances and post-training results on various auxiliary tasks, along with downstream imitation learning outcomes, are presented in Table 9.
@@
Model Diamond Sword Ladder Cooked Beef Iron Ingot
Qwen2-VL(raw) 0.53 0.40 0.03 0.10
Qwen2-VL-7B (one-stage) 0.10 0.40 0.07 0.13
ActVLP-Qwen2-VL 0.83 0.63 0.77 0.70
Table 9: Success rates comparing different training paradigms on selected Minecraft tasks. ’raw’ refers to the Qwen2-VL baseline fine-tuned directly. ’RT-2 (1-stage co-finetuning)’ corresponds to the Qwen2-VL-7B (one-stage) baseline. ’Act-VLP (3-stage post-training)’ represents our ActVLP-Qwen2-VL approach.